OpenStudy (anonymous):

Write 52 243 as a Decimal number - 1000

jimthompson5910 (jim_thompson5910):

\[5\frac{243}{1000}=5+\frac{243}{1000}=5+0.243=5.243\] So \[5\frac{243}{1000}=5.243\]

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

But gave you a metel anyway

jimthompson5910 (jim_thompson5910):

Oh sry, made a typo \[52\frac{243}{1000}=52.243\]